OBSERVATION OF BLUE-DETUNED PHOTOASSOCIATION TO THE 2 (0_{g}^{+}) STATE OF ^{85}Rb_{2} VIA REMPI

M. A. BELLOS, \underline{R. CAROLLO}, D. RAHMLOW, J. BANERJEE, M. BERMUDEZ, E. E. EYLER, P. L. GOULD, and W. C. STWALLEY, Department of Physics, University of Connecticut, Storrs, CT 06269.

We report photoassociation of ultracold atoms to vibrational levels blue of the ^{85}Rb_{2} 5s + 5p_{1/2} asymptote, in the previously-unobserved 2 (0_{g}^{+}) Hund’s case (c) state that corresponds to the 2 ^{1}Σ_{g}^{+} state in Hund’s case (a). These excitedstate ultracold molecules decay to the a^{3}Σ_{u}^{+} state and are detected by pulsed REMPI through the 2 ^{3}Σ_{g}^{+} state. We also see an order of magnitude enhancement in the v′ = 111, J′ = 5 ro-vibrational level of the 2 (0_{g}^{+}) state and present evidence for resonant coupling between this level and the v′ = 155 of the 2 (1_{g}) state, seen earlier in photoassociative trap loss^{a}. Following photoassociation to the observed levels of the 2 (0_{g}^{+}) state, spontaneous decay populates vibrational levels approximately halfway up the a^{3}Σ_{u}^{+} potential well, including levels v′′ = 18 though v′′ = 24. This pathway complements the blue-detuned photoassociation technique previously developed by this group^{b}, which accesses the bottom of the a state potential. This work is supported by the NSF and AFOSR.

aR. A. Cline, J. D. Miller, and D. J. Heinzen, Phys. Rev. Lett. 73, 632 (1994)

bM. A. Bellos, et. al., Phys. Chem. Chem. Phys. 13, 18880 (2011)